OpenStudy (anonymous):

simplify (-3x^4/y^2)^3

OpenStudy (anonymous):

you want formula or answer

OpenStudy (anonymous):

simplify/remove parenthesis

OpenStudy (anonymous):

-3^3 * x^12 / y^6 = -9x^(12) / y^6

OpenStudy (anonymous):

\[\Huge \frac{x^a}{x^b}=x^{a-b}\] \[\Huge \left(\frac{x}{y}\right)^c=\frac{x^c}{y^c}\]

OpenStudy (anonymous):

x^a/x^b=x^a-b

OpenStudy (anonymous):

and also \[\Huge \left(\frac{x^a}{y^b}\right)^c=\frac{x^{ac}}{y^{bc}}\]